# Optimal Rates for Generalization of Gradient Descent for Deep ReLU Classification

### v1

@@ A.2 Properties of Perturbation Terms
Let $\widehat{\mathbf{G}}^{a}_{b}(\mathbf{x})$ with $1\leq a\leq b\leq L$ be the matrix defined in (A.2). Assume $\max_{l\in[L]}\|\widehat{\mathbf{W}}^{l}\|_{2}\leq R\lesssim\sqrt{m}/(L^{2}\sqrt{\log m}),m\gtrsim L^{2}\log(nL/\delta)$ and $\widehat{\Sigma}^{l}(\mathbf{x}_{i})\in[-1,1]^{m\times m},\|\widehat{\Sigma}^{l}(\mathbf{x}_{i})\|_{0}\leq s\lesssim{m}/(L^{2}\log m)$ for all $i\in[n],l\in[L]$. Then, with probability at least $1-\delta$ for all $1\leq a\leq b\leq L,i\in[n]$, there holds
$\left\|\widehat{\mathbf{G}}^{a}_{b}(\mathbf{x}_{i})\right\|_{2}\lesssim L\sqrt{\frac{\log m}{m}}.$
@@
The proof is similar to that of Lemma 8.6 in Allen-Zhu et al. (2019), we give the proof here for the sake of completeness. We first prove that for any $1\leq a\leq b\leq L$,
$\left\|\sqrt{\frac{2}{m}}(\Sigma^{b}_{0}(\mathbf{x}_{i})+\widehat{\Sigma}^{b}(\mathbf{x}_{i}))\mathbf{W}^{b}(0)\cdots\sqrt{\frac{2}{m}}(\Sigma^{a+1}_{0}(\mathbf{x}_{i})+\widehat{\Sigma}^{a+1}(\mathbf{x}_{i}))\mathbf{W}^{a+1}(0)\right\|_{2}\lesssim L\sqrt{\log m}.$ (30)
We define a diagonal matrix $(\widehat{\Sigma}^{l}_{1}(\mathbf{x}_{i}))_{k,k}=\mathbb{I}\{\widehat{\Sigma}^{l}(\mathbf{x}_{i})_{k,k}\neq 0\}$, and $\|\widehat{\Sigma}^{l}_{1}(\mathbf{x}_{i})\|_{0}\leq s$. Therefore, $\widehat{\Sigma}^{l}(\mathbf{x}_{i})=\widehat{\Sigma}^{l}_{1}(\mathbf{x}_{i})\widehat{\Sigma}^{l}(\mathbf{x}_{i})\widehat{\Sigma}^{l}_{1}(\mathbf{x}_{i})$. We decompose the left term of (30) into $2^{b-a}$ terms and control them respectively. Each matrix can be written as (ignoring the superscripts and $\mathbf{x}_{i}$).
@@
$\left\|\sqrt{\frac{2}{m}}(\Sigma^{b}_{0}(\mathbf{x}_{i})+\widehat{\Sigma}^{b}(\mathbf{x}_{i}))\mathbf{W}^{b}(0)\cdots\sqrt{\frac{2}{m}}(\Sigma^{a+1}_{0}(\mathbf{x}_{i})+\widehat{\Sigma}^{a+1}(\mathbf{x}_{i}))\mathbf{W}^{a+1}(0)\right\|_{2}$
@@
Lemma 13 (Claim 11.2 and Proposition 11.3 in Allen-Zhu et al. (2019)).
@@
Lemma 14 (Theorem 2.8.4 in Vershynin (2018)).
@@
$\|h^{l}(\mathbf{x}_{i})-h^{l}_{0}(\mathbf{x}_{i})\|_{2}\lesssim L^{2}\sqrt{\frac{\log m}{m}}R\ \text{and}\ \|\Sigma^{l}(\mathbf{x}_{i})-\Sigma^{l}_{0}(\mathbf{x}_{i})\|_{0}\lesssim L^{4/3}(\log m)^{1/3}(mR)^{2/3}.$ (31)
@@
We prove these two inequalities by induction. Note that (31) holds for $l=0$. Now we suppose (31) holds for l1l-1. Let $\kappa>0$ be a constant. For $i\in[n]$ and $l\in[L]$, we define $A^{l}(\mathbf{x}_{i})=\{r\in[m]:\mathbb{I}\{\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}_{i})\rangle\geq 0\}\neq\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle\geq 0\}\}$, then $\|\Sigma^{l}(x_{i})-\Sigma^{l}_{0}(x_{i})\|_{0}=|A^{l}(\mathbf{x}_{i})|$. Furthermore, we decompose $A^{l}(\mathbf{x}_{i})$ into two parts based on the behavior of $\mathbf{w}^{l}_{r}(0)$:
$A^{l}_{1}(\mathbf{x}_{i})=\{r\in A^{l}(\mathbf{x}_{i}):|\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}_{i})\rangle|\leq\kappa\}\quad\text{and}\quad A_{2}^{l}(\mathbf{x}_{i})=\{r\in A^{l}(\mathbf{x}_{i}):|\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}_{i})\rangle|>\kappa\}.$
We will control $|A^{l}_{1}(\mathbf{x}_{i})|$ and $|A^{l}_{2}(\mathbf{x}_{i})|$ respectively.
For $r\in[m]$, we define $F_{r,i}^{l}=\mathbb{I}\{|\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}_{i})\rangle|\leq\kappa\}$. From Lemma 6 we know that $2/3\leq\|h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}^{2}\leq 4/3$. Condition on $h^{l-1}_{0}(\mathbf{x}_{i})$, $\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}_{i})\rangle\sim\mathcal{N}(0,\|h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}^{2})$, then we have
$\text{var}(F_{r,i}^{l})\leq\mathbb{E}(F_{r,i}^{l})^{2}=\mathbb{E}(F_{r,i}^{l})=\mathbb{P}(-\kappa\leq\langle\mathbf{W}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}_{i})\rangle\leq\kappa)$
@@
$\mathbb{P}\left(\left|\sum_{r=1}^{m}F_{r,i}^{l}-m\mathbb{E}(F_{r,i}^{l})\right|\leq mC\kappa|h^{l-1}_{0}(\mathbf{x}_{i})\right)\geq 1-2\exp\left(-\frac{(mC\kappa)^{2}/2}{mC\kappa+mC\kappa/3}\right).$
@@
$|A_{1}^{l}(\mathbf{x}_{i})|\leq\sum_{r=1}^{m}F_{r,i}^{l}\lesssim m\kappa.$ (32)
For $r\in A_{2}^{l}(\mathbf{x}_{i})$, since $\mathbb{I}\{\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}_{i})\rangle\geq 0\}\neq\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle\geq 0\}$, we have
$(\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}_{i})\rangle-\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)^{2}\geq|\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle|^{2}>\kappa^{2}.$
@@
$\|\mathbf{W}^{l}h^{l-1}(\mathbf{x}_{i})-\mathbf{W}^{l}(0)h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}^{2}\geq$ $\sum_{r\in A_{2}^{l}(\mathbf{x}_{i})}(\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}_{i})\rangle-\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)^{2}$
$>$ $\sum_{r\in A_{2}^{l}(\mathbf{x}_{i})}\kappa^{2}=\kappa^{2}|A_{2}^{l}(\mathbf{x}_{i})|.$ (33)
By assumption $\|h^{l-1}(\mathbf{x}_{i})-h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}\lesssim L^{2}R\sqrt{\log m/m}$ and Lemma 6, we get
$\|\mathbf{W}^{l}h^{l-1}(\mathbf{x}_{i})-\mathbf{W}^{l}(0)h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}^{2}$
$\leq$ $(\|\mathbf{W}^{l}-\mathbf{W}^{l}(0)\|_{2}\|h^{l-1}(\mathbf{x}_{i})-h^{l-1}_{0}(\mathbf{x}_{i})+h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}+\|\mathbf{W}^{l}(0)\|_{2}\|h^{l-1}(\mathbf{x}_{i})-h^{l-1}_{0}(\mathbf{x}_{i})\|_{2})^{2}$
$\lesssim$ $(R(\|h^{l-1}(\mathbf{x}_{i})-h^{l-1}_{0}(\mathbf{x}_{i})\|_{2}+1)+\sqrt{m}\|h^{l-1}(\mathbf{x}_{i})-h^{l-1}_{0}(\mathbf{x}_{i})\|_{2})^{2}\lesssim L^{4}R^{2}\log m.$
@@
$|A_{2}^{l}(\mathbf{x}_{i})|\lesssim\frac{L^{4}R^{2}\log m}{\kappa^{2}}.$ (34)
@@
$\|\Sigma^{l}(\mathbf{x}_{i})-\Sigma^{l}_{0}(\mathbf{x}_{i})\|_{0}=$ $|A^{l}(\mathbf{x}_{i})|=|A_{1}^{l}(\mathbf{x}_{i})|+|A_{2}^{l}(\mathbf{x}_{i})|$
@@
$\|\Sigma^{l}(\mathbf{x}_{i})-\Sigma^{l}_{0}(\mathbf{x}_{i})\|_{0}\lesssim L^{4/3}(\log m)^{1/3}(mR)^{2/3}\lesssim\frac{m}{L^{2}\log m}.$
@@
$h^{l}(\mathbf{x}_{i})-h^{l}_{0}(\mathbf{x}_{i})=\sum_{k=1}^{l}\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}_{i})(\mathbf{W}^{k}-\mathbf{W}^{k}(0))h_{0}^{k-1}(\mathbf{x}_{i}),$ (35)
where $\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}_{i})$ is defined as
$\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}_{i})=\left[\prod_{j=k+1}^{l}\sqrt{\frac{2}{m}}(\Sigma^{j}(\mathbf{x}_{i})+(\Sigma^{\prime\prime})^{j})\mathbf{W}^{j}\right]\sqrt{\frac{2}{m}}(\Sigma^{k}(\mathbf{x}_{i})+(\Sigma^{\prime\prime})^{k}).$ (36)
@@
$\|\Sigma^{j}(\mathbf{x}_{i})+(\Sigma^{\prime\prime})^{j}-\Sigma_{0}^{j}(\mathbf{x}_{i})\|_{0}$
$\leq$ $\|\Sigma^{j}(\mathbf{x}_{i})-\Sigma_{0}^{j}(\mathbf{x}_{i})\|_{0}+\|(\Sigma^{\prime\prime})^{j}\|_{0}\leq 2\|\Sigma^{j}(\mathbf{x}_{i})-\Sigma_{0}^{j}(\mathbf{x}_{i})\|_{0}\lesssim\frac{m}{L^{2}\log m}.$
@@
$\|\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}_{i})\|_{2}\lesssim L\sqrt{\frac{\log m}{m}}.$ (37)
@@
$\|h^{l}(\mathbf{x}_{i})-h^{l}_{0}(\mathbf{x}_{i})\|_{2}$
$=$ $\left\|\sum_{k=1}^{l}\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}_{i})(\mathbf{W}^{k}-\mathbf{W}^{k}(0))h_{0}^{k-1}(\mathbf{x}_{i})\right\|_{2}$
$\lesssim$ $\sum_{k=1}^{l}L\sqrt{\frac{\log m}{m}}R\|h_{0}^{k-1}(\mathbf{x}_{i})\|_{2}\lesssim L^{2}\sqrt{\frac{\log m}{m}}R,$
@@
$\|h^{l}(\mathbf{x}_{i})\|_{2}\lesssim L^{2}\sqrt{\frac{\log m}{m}}R+1\lesssim 1.$ (38)
@@
Although our approach shares similarities with Lemma B.3 in Zou et al. (2018), our analysis significantly relaxes the required conditions. Specifically, we only require $R/\sqrt{m}=\widetilde{O}(L^{-5})$, whereas their result demands the stricter scaling $R/\sqrt{m}=\widetilde{O}(L^{-11})$.
Furthermore, compared to Lemma 8.2 in Allen-Zhu et al. (2019), they derive the bound $\|h^{l}(\mathbf{x}_{i})-h^{l}_{0}(\mathbf{x}_{i})\|_{2}\lesssim RL^{5/2}\sqrt{\log m}/\sqrt{m}$, which is worse than our result by a factor of $\sqrt{L}$.
@@
We consider the $1/(C^{L}\sqrt{m})$-cover of $S^{d-1}$ and denote it by $D=\{\mathbf{x}^{1},\cdots,\mathbf{x}^{|D|}\}$. By Lemma 4.2.13 in Vershynin (2018),
@@
Note that Lemma 15 holds for $n$ points $\mathbf{x}_{1},\cdots,\mathbf{x}_{n}$. We consider $|D|$ points in $D$, and obtain that if $m\gtrsim L^{11}d(\log m)^{5}\log(L/\delta)R^{2}$, then $m\gtrsim L^{10}\log(|D|L/\delta)(\log m)^{4}R^{2}$. Hence,
@@
Remark 9. Compared to prior work, while Allen-Zhu et al. (2019); Zou et al. (2018) only establishes bounds for the training data, we prove the uniform convergence over the entire input space. Previous work on uniform concentration demonstrated that $\sup_{\mathbf{x}\in\mathcal{X}}\|h^{l}(\mathbf{x})-h_{0}^{l}(\mathbf{x})\|_{2}\lesssim C^{L}R/\sqrt{m}$ (Xu and Zhu, 2024). We present a significant improvement, reducing the dependence on $L$ from exponential to polynomial.

### v2

@@ A.2 Properties of Perturbation Terms
Let $\widehat{\mathbf{G}}^{a}_{b}(\mathbf{x})$ with $1\leq a\leq b\leq L$ be the matrix defined in (A.2). Assume $\max_{l\in[L]}\|\widehat{\mathbf{W}}^{l}\|_{2}\leq R\lesssim\sqrt{m}/(L^{2}\sqrt{\log m}),m\gtrsim L^{2}\log(NL/\delta)$ and $\widehat{\Sigma}^{l}(\mathbf{x}^{i})\in[-1,1]^{m\times m},\|\widehat{\Sigma}^{l}(\mathbf{x}^{i})\|_{0}\leq s\lesssim{m}/(L^{2}\log m)$ for all $i\in[N],l\in[L]$. Then, with probability at least $1-\delta$ for all $1\leq a\leq b\leq L,i\in[N]$, there holds
$\left\|\widehat{\mathbf{G}}^{a}_{b}(\mathbf{x}^{i})\right\|_{2}\lesssim L\sqrt{\frac{\log m}{m}}.$
@@
The proof is similar to that of Lemma 8.6 in pmlr-v97-allen-zhu19a, we give the proof here for the sake of completeness. We first prove that for any $1\leq a\leq b\leq L$,
$\left\|\sqrt{\frac{2}{m}}(\Sigma^{b}_{0}(\mathbf{x}^{i})+\widehat{\Sigma}^{b}(\mathbf{x}^{i}))\mathbf{W}^{b}(0)\cdots\sqrt{\frac{2}{m}}(\Sigma^{a+1}_{0}(\mathbf{x}^{i})+\widehat{\Sigma}^{a+1}(\mathbf{x}^{i}))\mathbf{W}^{a+1}(0)\right\|_{2}\lesssim L\sqrt{\log m}.$ (30)
We define a diagonal matrix $(\widehat{\Sigma}^{l}_{1}(\mathbf{x}^{i}))_{k,k}=\mathbb{I}\{\widehat{\Sigma}^{l}(\mathbf{x}^{i})_{k,k}\neq 0\}$, and $\|\widehat{\Sigma}^{l}_{1}(\mathbf{x}^{i})\|_{0}\leq s$. Therefore, $\widehat{\Sigma}^{l}(\mathbf{x}^{i})=\widehat{\Sigma}^{l}_{1}(\mathbf{x}^{i})\widehat{\Sigma}^{l}(\mathbf{x}^{i})\widehat{\Sigma}^{l}_{1}(\mathbf{x}^{i})$. We decompose the left term of (30) into $2^{b-a}$ terms and control them respectively. Each matrix can be written as (ignoring the superscripts and $\mathbf{x}^{i}$).
@@
$\left\|\sqrt{\frac{2}{m}}(\Sigma^{b}_{0}(\mathbf{x}^{i})+\widehat{\Sigma}^{b}(\mathbf{x}^{i}))\mathbf{W}^{b}(0)\cdots\sqrt{\frac{2}{m}}(\Sigma^{a+1}_{0}(\mathbf{x}^{i})+\widehat{\Sigma}^{a+1}(\mathbf{x}^{i}))\mathbf{W}^{a+1}(0)\right\|_{2}$
@@
Lemma 13 (Claim 11.2 and Proposition 11.3 in pmlr-v97-allen-zhu19a).
@@
Lemma 14 (Theorem 2.8.4 in vershynin2018high).
@@
$\|h^{l}(\mathbf{x}^{i})-h^{l}_{0}(\mathbf{x}^{i})\|_{2}\lesssim L^{2}\sqrt{\frac{\log m}{m}}R\ \text{and}\ \|\Sigma^{l}(\mathbf{x}^{i})-\Sigma^{l}_{0}(\mathbf{x}^{i})\|_{0}\lesssim L^{4/3}(\log m)^{1/3}(mR)^{2/3}.$ (31)
@@
We prove these two inequalities by induction. Note that (31) holds for $l=0$. Now we suppose (31) holds for l1l-1. Let $\kappa>0$ be a constant. For $i\in[N]$ and $l\in[L]$, we define $A^{l}(\mathbf{x}^{i})=\{r\in[m]:\mathbb{I}\{\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}^{i})\rangle\geq 0\}\neq\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle\geq 0\}\}$, then $\|\Sigma^{l}(\mathbf{x}^{i})-\Sigma^{l}_{0}(\mathbf{x}^{i})\|_{0}=|A^{l}(\mathbf{x}^{i})|$. Furthermore, we decompose $A^{l}(\mathbf{x}^{i})$ into two parts based on the behavior of $\mathbf{w}^{l}_{r}(0)$:
$A^{l}_{1}(\mathbf{x}^{i})=\{r\in A^{l}(\mathbf{x}^{i}):|\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}^{i})\rangle|\leq\kappa\}\quad\text{and}\quad A_{2}^{l}(\mathbf{x}^{i})=\{r\in A^{l}(\mathbf{x}^{i}):|\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}^{i})\rangle|>\kappa\}.$
We will control $|A^{l}_{1}(\mathbf{x}^{i})|$ and $|A^{l}_{2}(\mathbf{x}^{i})|$ respectively.
For $r\in[m]$, we define $F_{r,i}^{l}=\mathbb{I}\{|\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}^{i})\rangle|\leq\kappa\}$. From Lemma 6 we know that $2/3\leq\|h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}^{2}\leq 4/3$. Condition on $h^{l-1}_{0}(\mathbf{x}^{i})$, $\langle\mathbf{w}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}^{i})\rangle\sim\mathcal{N}(0,\|h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}^{2})$, then we have
$\text{var}(F_{r,i}^{l})\leq\mathbb{E}(F_{r,i}^{l})^{2}=\mathbb{E}(F_{r,i}^{l})=\mathbb{P}(-\kappa\leq\langle\mathbf{W}_{r}^{l}(0),h^{l-1}_{0}(\mathbf{x}^{i})\rangle\leq\kappa)$
@@
$\mathbb{P}\left(\left|\sum_{r=1}^{m}F_{r,i}^{l}-m\mathbb{E}(F_{r,i}^{l})\right|\leq mC\kappa|h^{l-1}_{0}(\mathbf{x}^{i})\right)\geq 1-2\exp\left(-\frac{(mC\kappa)^{2}/2}{mC\kappa+mC\kappa/3}\right).$
@@
$|A_{1}^{l}(\mathbf{x}^{i})|\leq\sum_{r=1}^{m}F_{r,i}^{l}\lesssim m\kappa.$ (32)
For $r\in A_{2}^{l}(\mathbf{x}_{i})$, since $\mathbb{I}\{\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}^{i})\rangle\geq 0\}\neq\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle\geq 0\}$, we have
$(\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}^{i})\rangle-\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)^{2}\geq|\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle|^{2}>\kappa^{2}.$
@@
$\|\mathbf{W}^{l}h^{l-1}(\mathbf{x}^{i})-\mathbf{W}^{l}(0)h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}^{2}\geq$ $\sum_{r\in A_{2}^{l}(\mathbf{x}_{i})}(\langle\mathbf{w}_{r}^{l},h^{l-1}(\mathbf{x}^{i})\rangle-\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)^{2}$
$>$ $\sum_{r\in A_{2}^{l}(\mathbf{x}^{i})}\kappa^{2}=\kappa^{2}|A_{2}^{l}(\mathbf{x}^{i})|.$ (33)
By assumption $\|h^{l-1}(\mathbf{x}^{i})-h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}\lesssim L^{2}R\sqrt{\log m/m}$ and Lemma 6, we get
$\|\mathbf{W}^{l}h^{l-1}(\mathbf{x}^{i})-\mathbf{W}^{l}(0)h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}^{2}$
$\leq$ $(\|\mathbf{W}^{l}-\mathbf{W}^{l}(0)\|_{2}\|h^{l-1}(\mathbf{x}^{i})-h^{l-1}_{0}(\mathbf{x}^{i})+h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}+\|\mathbf{W}^{l}(0)\|_{2}\|h^{l-1}(\mathbf{x}^{i})-h^{l-1}_{0}(\mathbf{x}^{i})\|_{2})^{2}$
$\lesssim$ $(R(\|h^{l-1}(\mathbf{x}^{i})-h^{l-1}_{0}(\mathbf{x}^{i})\|_{2}+1)+\sqrt{m}\|h^{l-1}(\mathbf{x}^{i})-h^{l-1}_{0}(\mathbf{x}^{i})\|_{2})^{2}\lesssim L^{4}R^{2}\log m.$
@@
$|A_{2}^{l}(\mathbf{x}^{i})|\lesssim\frac{L^{4}R^{2}\log m}{\kappa^{2}}.$ (34)
@@
$\|\Sigma^{l}(\mathbf{x}^{i})-\Sigma^{l}_{0}(\mathbf{x}^{i})\|_{0}=$ $|A^{l}(\mathbf{x}^{i})|=|A_{1}^{l}(\mathbf{x}^{i})|+|A_{2}^{l}(\mathbf{x}^{i})|$
@@
$\|\Sigma^{l}(\mathbf{x}^{i})-\Sigma^{l}_{0}(\mathbf{x}^{i})\|_{0}\lesssim L^{4/3}(\log m)^{1/3}(mR)^{2/3}\lesssim\frac{m}{L^{2}\log m}.$
@@
$h^{l}(\mathbf{x}^{i})-h^{l}_{0}(\mathbf{x}^{i})=\sum_{k=1}^{l}\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}^{i})(\mathbf{W}^{k}-\mathbf{W}^{k}(0))h_{0}^{k-1}(\mathbf{x}^{i}),$ (35)
where $\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}^{i})$ is defined as
$\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}^{i})=\left[\prod_{j=k+1}^{l}\sqrt{\frac{2}{m}}(\Sigma^{j}(\mathbf{x}^{i})+(\Sigma^{\prime\prime})^{j})\mathbf{W}^{j}\right]\sqrt{\frac{2}{m}}(\Sigma^{k}(\mathbf{x}^{i})+(\Sigma^{\prime\prime})^{k}).$ (36)
@@
$\|\Sigma^{j}(\mathbf{x}^{i})+(\Sigma^{\prime\prime})^{j}-\Sigma_{0}^{j}(\mathbf{x}^{i})\|_{0}$
$\leq$ $\|\Sigma^{j}(\mathbf{x}^{i})-\Sigma_{0}^{j}(\mathbf{x}^{i})\|_{0}+\|(\Sigma^{\prime\prime})^{j}\|_{0}\leq 2\|\Sigma^{j}(\mathbf{x}^{i})-\Sigma_{0}^{j}(\mathbf{x}^{i})\|_{0}\lesssim\frac{m}{L^{2}\log m}.$
@@
$\|\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}^{i})\|_{2}\lesssim L\sqrt{\frac{\log m}{m}}.$ (37)
@@
$\|h^{l}(\mathbf{x}^{i})-h^{l}_{0}(\mathbf{x}^{i})\|_{2}$
$=$ $\left\|\sum_{k=1}^{l}\widehat{\mathbf{G}}_{l,0}^{k}(\mathbf{x}^{i})(\mathbf{W}^{k}-\mathbf{W}^{k}(0))h_{0}^{k-1}(\mathbf{x}^{i})\right\|_{2}$
$\lesssim$ $\sum_{k=1}^{l}L\sqrt{\frac{\log m}{m}}R\|h_{0}^{k-1}(\mathbf{x}^{i})\|_{2}\lesssim L^{2}\sqrt{\frac{\log m}{m}}R,$
@@
$\|h^{l}(\mathbf{x}^{i})\|_{2}\lesssim L^{2}\sqrt{\frac{\log m}{m}}R+1\lesssim 1.$ (38)
@@
Although our approach shares similarities with Lemma B.3 in zou2018stochasticgradientdescentoptimizes, our analysis significantly relaxes the required conditions. Specifically, we only require $R/\sqrt{m}=\widetilde{O}(L^{-5})$, whereas their result demands the stricter scaling $R/\sqrt{m}=\widetilde{O}(L^{-11})$.
Furthermore, compared to Lemma 8.2 in pmlr-v97-allen-zhu19a, they derive the bound $\|h^{l}(\mathbf{x}^{i})-h^{l}_{0}(\mathbf{x}^{i})\|_{2}\lesssim RL^{5/2}\sqrt{\log m}/\sqrt{m}$, which is worse than our result by a factor of $\sqrt{L}$.
@@
We consider the $1/(C^{L}\sqrt{m})$-cover of $S^{d-1}$ and denote it by $D=\{\mathbf{x}^{1},\cdots,\mathbf{x}^{|D|}\}$. By Lemma 4.2.13 in vershynin2018high,
@@
Note that Lemma 15 holds for any finite set $K=\{\mathbf{x}^{1},\cdots,\mathbf{x}^{N}\}$. Letting $K=D$, we obtain that if $m\gtrsim L^{11}d(\log m)^{5}\log(L/\delta)R^{2}\gtrsim L^{10}\log(|D|L/\delta)(\log m)^{4}R^{2}$, then
@@
Remark 9. This lemma is a property of deep ReLU networks near initialization that does not depend on the training data. Compared to prior work, while pmlr-v97-allen-zhu19a; zou2018stochasticgradientdescentoptimizes only establishes bounds for the training data, we prove the uniform convergence over the entire input space. Previous work on uniform concentration demonstrated that $\sup_{\mathbf{x}\in\mathcal{X}}\|h^{l}(\mathbf{x})-h_{0}^{l}(\mathbf{x})\|_{2}\lesssim C^{L}R/\sqrt{m}$ (JMLR:v25:23-0740). We present a significant improvement, reducing the dependence on $L$ from exponential to polynomial.
In the following part, we apply previous technical lemmas to $K=S_{1}$ and get properties of deep neural networks over the training dataset.